Operation Summary for August 27, 2010 - August 30, 2010 |
The Integrated Luminosity for the period from 8/23/10 to 8/30/10 was 41.23 pb-1. Friday August 27 The day shift began with Operations monitoring the Tevatron (TeV) store 8023 (57.07E30) and the Recycler stash (237.7E10), with the Antiproton Source (Pbar) stacking (30.7E10), with the Linac, Booster and the Main Injector (MI) operating normally, with the MiniBooNE and NuMI experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, Operations terminated store 8023 at 11 AM, put the TeV into HEP shot setup, and established store 8024 at 12:30 PM with an average initial luminosity (AIL) of 282.4E30. RF personnel investigated and replaced the Recycler amplifier KW4. During the evening shift, Operations monitored beam to Booster, MI, Pbar, NuMI, MiniBooNE, and when available, transferred antiprotons to the Recycler. Saturday August 28 The midnight shift began with Operations monitoring store 8024 (82.2E30) and the stash (279.8E10), with Pbar stacking (14.8E10), and with MiniBooNE taking beam. During the midnight shift, Operations terminated store 8024 at 2 AM, put the TeV into shot setup, and established store 8025 at 3:20 AM with an AIL of 278.5E30. During the day shift, Operations terminated store 8025 at noon and then prepared D0 and CDF for an access. Operators checked the TeV orbits and then put the TeV into shot setup at 4:15 PM. The final antiproton transfer into the TeV was lost when the MI-30 interlocks dropped out due to a PLC microprocessor failure. Then, unrelated to the interlock problem, the D0 low beta magnets quenched prior to initiating collisions. Investigations found that the critical angle bumps at the interaction points (IP) were inadvertently removed causing bad beam positions and a subsequent quench. Operators contacted interlock personnel to troubleshoot and replace the PLC at MI-30. Following the repairs, operators search and secured MI-30. Prior to reestablishing the MI ramp, the NuMI extraction kicker tripped off on a low fluorinert liquid level indication that required another access to MI-30. Operators called in Mechanical Support personnel to assist. During the evening shift, Operations reported that the shift began with efforts to recover the TeV from the D0 low beta quench and with filling the NuMI kicker with fluorinert. Operations completed the TeV checkout at 9:30 PM. Mechanical Support personnel came out of the MI access at 9:30 PM and Operations reestablished MI and NuMI beam at 10 PM. Sunday August 29 The midnight shift began with the TeV being checked out, with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations put the TeV into shot setup at 4:30 AM and established store 8034 at 5:35 AM with an AIL of 168E30. A Linac RF power amplifier (LRF2) breaker tripped off at 6:15 PM; operators called in Linac personnel. During the day shift, Linac personnel addressed the auxiliary breaker problem with LRF2 and a low-level RF problem with LRF1. Linac resumed sending beam at 10 AM. During the evening shift, Operations terminated store 8034 at 9:45 AM and put the TeV into shot setup. Unfortunately, the setup ended abruptly with a quench at TeV sectors A3 and A1. Following the quench, the B3 power supply indicated an electronic dump failure. Operators called in EE Support personnel to investigate; they found and replaced a failed IGCT (integrated gate-commutated thyristor). Monday August 30 The midnight shift began with the TeV recovering from a quench, with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, EE Support personnel reported finding no obvious beam-related causes for the quench and indicated that the failed IGCT was most likely a symptom of the B3 power supply going into bypass. Operators checked out the B3 dump and then had the TeV ramping by 5 AM. Operations checked the orbits. Operators contacted TeV personnel to verify that the appropriate IP angle bumps were included. The Plans for Monday The plans for today are to stack, stash, and store. The Run Coordinator will allow MI and Recycler studies. |
|
More Information Comments and Suggestions |
| last modified 08/30/2010 email Fermilab |
| Security, Privacy, Legal | |